OpenStudy (anonymous):

The area of a circle is given by: A(r) = πr squared, where A is the area and r is the radius of a circle. Express the radius r as a function of the Area.

OpenStudy (anonymous):

r(A) = 2Aπ ?

ganeshie8 (ganeshie8):

i think you need to divide and take square root

ganeshie8 (ganeshie8):

\[\large A = \pi r^2\]

ganeshie8 (ganeshie8):

to express radius as a function of something, we need to isolate \(r\)

ganeshie8 (ganeshie8):

start by dividing \(\pi\) both sides

OpenStudy (anonymous):

A= 4

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

this is just algebra gimmick, you don't need to *actually* divide stuff.. just put the \(\pi\) under A, thats all...

ganeshie8 (ganeshie8):

\[\large A = \pi r^2\] divide \(\pi \) both sides, you get : \[\large \dfrac{A}{\pi} = r^2\]

ganeshie8 (ganeshie8):

take sqrt both sides, you get : \[\large \pm \sqrt{\dfrac{A}{\pi}} = r\]

ganeshie8 (ganeshie8):

since radius is defined as the `distance` between center and point on circle, it can never be negative. so discard the negative value

ganeshie8 (ganeshie8):

\[\large r = \sqrt{\dfrac{A}{\pi}}\]

ganeshie8 (ganeshie8):

Oh yes, since it depends on A, we say it is a function of A, and write it in function notation like below : \[\large r(A) = \sqrt{\dfrac{A}{\pi}}\]

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

we're done !

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

a or c was it?

OpenStudy (anonymous):

I dont wanna get it wrong so just checing

OpenStudy (anonymous):

checking

OpenStudy (anonymous):

@Cosmichaotic

ganeshie8 (ganeshie8):

are you giving an online exam ?

OpenStudy (anonymous):

no its a practice

OpenStudy (anonymous):

quiz

ganeshie8 (ganeshie8):

fine then :)

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

wait so a or c?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

clearly i have failed in explaining properly >.< tell me why do you it is a or c ?

OpenStudy (anonymous):

because your dividing

OpenStudy (anonymous):

so its pie over A I know that

ganeshie8 (ganeshie8):

You have : \(\large A = \pi r^2\)

ganeshie8 (ganeshie8):

divide \(\pi\) both sides, what do you get ?

OpenStudy (anonymous):

uhh

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

then

OpenStudy (anonymous):

thats my best guess

ganeshie8 (ganeshie8):

I'm not sure if you're able to see the latex code on your end...

ganeshie8 (ganeshie8):

if the answer helps - the correct answer is the last option

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pellet

OpenStudy (anonymous):

I got a 42 percent

OpenStudy (anonymous):

@ganeshie8